# Shoto Aoki       Mikio Furuta    Shinichiroh Matsuo    Tetsuya Onogi    Satoshi Yamaguchi

### v1

@@ 1 Introduction
where $\eta(H_{W})$ is known as the $\eta$ invariant in mathematics, which was introduced by Atiyah, Patodi and Singer [10]. In this talk, we show a deeper mathematical meaning of the right-hand side of the equality, and try to convince the readers that the massive Wilson Dirac operator is an equally good or even better object in terms of $K$-theory [11, 12] than the overlap Dirac operator to describe the gauge field topology111 There are some developments from mathematicians. The Atiyah-Singer index theorem was directly formulated in [13] on a lattice using the algebraic index theorem of Nest and Tsygan. In [14], the lattice approximation of analytic indices through the higher index theory of almost flat vector bundles was formulated. .
@@ 2 $K$-theory
$K$-theory [11, 12] is one of the generalized cohomology theories which is useful to classify the vector bundles. Here we briefly describe the essence of $K$-theory and how it describes the index of the Dirac operators.
@@
The element of $K^{0}(X)$ group is given by a set $[E_{1},E_{2}]$ where $E_{1}$ and $E_{2}$ are two vector bundles over the same base space $X$. Equivalently, we can consider an operator and its conjugate,
@@
where $\mathcal{H}_{E}$ is the whole Hilbert space on which $D$ acts222 This map is analogous to the forgetful map from a complex number to two real numbers $f:\mathbb{C}\to\mathbb{R}^{2}$ where the map is given by $a+ib\to(a,b)$ or the type conversion in program codes from complicated one to simple one like from “str” to “list”. . Many information is lost with the map, but one (the Dirac operator index) remains.
In this work, we would like to suspend the “point” to an interval $I$ with two ends denoted by $\partial I$. In $K$-theory, there is an important isomorphism called the suspension isomorphism $K^{0}(\mbox{point})\cong K^{1}(I,\partial I)$ where the one-to-one map
@@
exists where the one-parameter family of the Dirac operator is labeled by $t\in I=[-1,1]$ with a condition $D_{\pm 1}$ is invertible. $p^{*}$ denotes the pullback of the projection $p:I\to\mbox{point}$. In the following analysis, we take $I$ as a parameter space of the fermion mass, and the same Hilbert space $\mathcal{H}_{E}$ is extended along $I$ in a trivial way. We, therefore, omit $p^{*}$ for simplicity in the following. The physical meaning of the isomorphism will be given soon later.
@@
In the standard formulation of the Dirac operator index, we need a massless Dirac operator and its zero modes with definite chirality. Therefore, we need to consider the $K^{0}(\mbox{point})$ elements given by $[\mathcal{H}_{E},D,\gamma]$. But we will show that it is isomorphic to $K^{1}(I,\partial I)$ whose element is described by massive Dirac operators $[p^{*}\mathcal{H}_{E},\gamma(D+m)]$ where the mass term is varied in the range $-M\leq m\leq+M$.
@@
Let us illustrate in Fig. 1, the spectrum of this massive Dirac operator as a function of mass parameter $m\in[-M,M]$. The nonzero modes make parabolic (dashed-green) curves which is $\pm$ symmetric. On the other hand, the chiral modes cross zero from negative to positive when its chirality is $+$ (red-solid line) and go from positive to negative when they have the negative chirality (blue-solid).
@@
This describes the physical meaning of the suspension isomorphism $K^{0}(\mbox{point})\cong K^{1}(I,\partial I)$. The standard definition of the Dirac operator index which characterizes $[\mathcal{H}_{E},D,\gamma]$ is always equal to the spectral flow along the one parameter family $-M\leq m\leq+M$ of the massive Dirac operator $\gamma(D+m)$ which characterizes $[p^{*}\mathcal{H}_{E},\gamma(D+m)]$. The massless definition corresponds to counting the index by points at $m=0$, while the massive case counts the same index by lines in the range $[-M,M]$. The two definitions of the index always agree.
@@
What happens when a chiral symmetry breaking regularization is employed (on a lattice)? The spectrum will be deformed like Fig. 2. The nonzero modes are no more $\pm$ symmetric and the chiral modes do not cross zero at $m=0$. It is difficult to identify the chiral zero modes to give the standard definition of the Dirac operator index. However, we can still count the crossing red and blue lines as far as the two end points at $m=\pm M$ have a enough gap from zero in the spectrum. Counting lines is easier and stabler against the symmetry breaking than counting points. This is the core of this work and explains why the Wilson Dirac operator works as the $K^{1}(I,\partial I)$ group element to describe the index on the lattice. The $K^{1}(I,\partial I)$ group is insensitive to the existence of the chirality operator $\gamma$ and easier to consider in lattice gauge theory.
@@ 4 Main theorem
We regularize the torus $T^{2n}$ by a square lattice with lattice spacing $a$. Note that the fiber vector space is kept continuous. There is some ambiguity in defining the link variables. The standard approach is to define the continuum gauge field first and then to define the link variables by the Wilson line, taking the path-ordered product of the exponentiated gauge field connection. In this work, we take an opposite direction to simplify the analysis. We first define the link variables between arbitrary two points $(x,y)$ on the torus:
@@
where $\nabla_{\mu}$ is the forward difference operator and $\nabla_{\mu}^{*}$ is its conjugate. Here we have chosen the Wilson coefficient unity.
@@
which is given by the difference $[(\mathcal{H}_{\text{cont.}},\gamma(D_{\rm cont.}+m))]-[(\mathcal{H}_{\text{lat.}},\gamma(D_{W}+m))]$ and adding a perturbation: the continuum-lattice mixing mass term $tf_{a}$ and its conjugate. Here, $f_{a}$ is a map from $\mathcal{H}_{\text{lat.}}$ to $\mathcal{H}_{\text{cont.}}$, interpolating the vectors on the discrete lattice sites to those on the continuum position space by a product of simple linear division of unity in every direction. The conjugate $f_{a}^{*}$ corresponds to the coarse graining map from the continuum Hilbert space to the discrete space on the lattice.
We prove by contradiction for any gauge field background that there exists a finite lattice spacing $a_{0}$, such that for any lattice spacing $a<a_{0}$, $\hat{D}$ is always invertible, having no zero mode, along the path in the two parameter space $(m,t)=(-M,0)\to(-M,1)\to(+M,1)\to(+M,0)$. This is a sufficient condition for $[(\mathcal{H}_{\text{cont.}},\gamma(D_{\rm cont.}+m))]$ and $[(\mathcal{H}_{\text{lat.}},\gamma(D_{W}+m))]$ to be in the same equivalence class. We can show that the $\eta(\gamma(D_{W}+M))=0$ for $M>0$ and therefore, we obtain
@@ 5 Comparison with the overlap Dirac index
One possible application is the index on manifolds with boundaries. In the previous sections, we have only considered a periodic square lattice, whose continuum limit is a flat torus. For the massive Dirac operators, it is straightforward to introduce a free boundary condition, to obtain the domain-wall fermion Dirac operator $D_{\rm DW}$ [22, 23, 24] In [15, 16] we have shown in continuum theory that the $\eta$ invariant of domain-wall Dirac operator is equal to the Atiyah-Patodi-Singer(APS) index of the Dirac operator with a nontrivial boundary. Therefore, it is natural to conjecture that $-\frac{1}{2}\eta(\gamma(D_{\rm DW}))$ at sufficiently small lattice spacings agrees with the APS index. The overlap Dirac operator, in contrast, is no more chiral symmetric when we impose such a boundary condition, since the boundary condition invalidates the Ginsparg-Wilson relation [25].
@@
We also expect that a nontrivial gravitational background can be implemented by the curved domain-wall mass term [27, 28, 29, 30, 31, 31, 32, 33]. It would be interesting to formulate, for instance, the $\hat{A}$ genus of the curved domain-wall on a lattice.

### v2

@@ 1 Introduction
where $\eta(H_{W})$ is known as the $\eta$ invariant in mathematics, which was introduced by Atiyah, Patodi and Singer [10]. In this talk, we show a $K$-theoretic meaning of the right-hand side of the equality, and try to convince the readers that the massive Wilson Dirac operator is an equally good or even better mathematical object in terms of $K$-theory [11, 12] than the overlap Dirac operator to describe the gauge field topology111 Recently there are some developments in mathematics. The Atiyah-Singer index theorem was directly formulated in [13] on a lattice using the algebraic index theorem of Nest and Tsygan. In [14], the lattice approximation of analytic indices through the higher index theory of almost flat vector bundles was formulated. .
@@ 2 $K$-theory
$K$-theory [11, 12] is one of the generalized cohomology theories and useful in classifying the vector bundles. Here we briefly describe the essence of $K$-theory and how it describes the index of the Dirac operators.
@@
The element of $K^{0}(X)$ group is given by a set $[E_{1},E_{2}]$ where $E_{1}$ and $E_{2}$ are two vector bundles over the same base space $X$. Equivalently, we can consider an operator and its conjugate222 To be precise, the operators act on the sections of $E_{i}$. ,
@@
where $\mathcal{H}_{E}$ is the whole Hilbert space on which $D$ acts333 This map is analogous to the forgetful map from a complex number to two real numbers $f:\mathbb{C}\to\mathbb{R}^{2}$ where the map is given by $a+ib\to(a,b)$ or the type conversion in program codes from complicated one to simple one like from “str” to “list”. . Many information is lost with the map, but one (the Dirac operator index) remains.
In this work, we would like to suspend the “point” to an interval $I$ with two ends denoted by $\partial I$. In $K$-theory, there is an important isomorphism called the suspension isomorphism $K^{0}(\mbox{point})\cong K^{1}(I,\partial I)$ with the one-to-one map
@@
exists where the one-parameter family of the Dirac operator $D_{t}$ is labeled by $t\in I=[-1,1]$ with a condition that $D_{\pm 1}$ is invertible. $p^{*}$ denotes the pullback of the projection $p:I\to\mbox{point}$. In the following analysis, we take $I$ as a parameter space of the fermion mass, and the same Hilbert space $\mathcal{H}_{E}$ is extended along $I$ in a trivial way. We, therefore, omit $p^{*}$ for simplicity in the following. The physical meaning of the isomorphism will be given soon later.
@@
In the standard formulation of the Dirac operator index, we need a massless Dirac operator and its zero modes with definite chirality. Therefore, we need to consider the $K^{0}(\mbox{point})$ elements given by $[\mathcal{H}_{E},D,\gamma]$. But we will show that it is isomorphic to $K^{1}(I,\partial I)$ whose element is described by the massive Dirac operators $[\mathcal{H}_{E},\gamma(D+m)]$ where the mass term is varied in the range $-M\leq m\leq+M$.
@@
Let us illustrate in Fig. 1 the spectrum of this massive Dirac operator as a function of mass parameter $m\in[-M,M]$. The nonzero modes make parabolic (dashed-green) curves which are $\pm$ symmetric. On the other hand, the chiral modes cross zero from negative to positive when its chirality is $+$ (red-solid line) and go from positive to negative when they have the negative chirality (blue-solid).
@@
This describes the physical meaning of the suspension isomorphism $K^{0}(\mbox{point})\cong K^{1}(I,\partial I)$. The standard definition of the Dirac operator index which characterizes $[\mathcal{H}_{E},D,\gamma]$ is always equal to the spectral flow along the one parameter family $-M\leq m\leq+M$ of the massive Dirac operator $\gamma(D+m)$ which characterizes $[p^{*}\mathcal{H}_{E},\gamma(D+m)]$. The massless definition corresponds to counting the index by points at $m=0$, while the massive case counts the same index by lines crossing zero in the range $[-M,M]$. The two definitions of the index always agree.
@@
What happens when a chiral symmetry breaking regularization is employed (on a lattice)? The spectrum will be deformed like Fig. 2. The nonzero modes are no more $\pm$ symmetric and the chiral modes do not cross zero at $m=0$. It is difficult to identify the chiral zero modes to give the standard definition of the Dirac operator index. However, we can still count the zero-crossing lines as far as the two end points at $m=\pm M$ have a enough gap from zero in the spectrum. Counting lines is easier and stabler against the symmetry breaking than counting points. This is the core of this work and explains why the Wilson Dirac operator works as the $K^{1}(I,\partial I)$ group element to describe the index on the lattice. The $K^{1}(I,\partial I)$ group is insensitive to the existence of the chirality operator $\gamma$ and easier to consider in lattice gauge theory.
@@ 4 Main theorem
We regularize the torus $T^{2n}$ by a square lattice with a lattice spacing $a$. Note that the fiber vector space is kept continuous. There is some ambiguity in defining the link variables from continuum theory. The standard approach is to define the continuum gauge field first and then to define the link variables by the Wilson line, taking the path-ordered product of the exponentiated gauge field connection. In this work, we take an opposite direction to simplify the analysis. We first define the link variables between arbitrary two points $(x,y)$ on the torus:
@@
where $\nabla_{\mu}$ is the forward covariant difference operator acting as $\nabla_{\mu}\psi(x)=[U(x,x+ae_{\mu})\psi(x+ae_{\mu})-\psi(x)]/a$, and $\nabla_{\mu}^{*}$ is its conjugate. Here we have chosen the Wilson coefficient unity.
@@
which is given by the difference $[(\mathcal{H}_{\text{cont.}},\gamma(D_{\rm cont.}+m))]-[(\mathcal{H}_{\text{lat.}},\gamma(D_{W}+m))]$ and adding a perturbation: the continuum-lattice mixing mass term $tf_{a}$ and its conjugate. Here, $f_{a}$ is a map from $\mathcal{H}_{\text{lat.}}$ to $\mathcal{H}_{\text{cont.}}$, interpolating the vectors on the discrete lattice sites to those on the continuum position space by a product of simple linear partition of unity in every direction. The conjugate $f_{a}^{*}$ corresponds to the coarse graining map from the continuum Hilbert space to the discrete space on the lattice.
We prove by contradiction for any gauge field background determined by $\{U(x,y)\}$ that there exists a finite lattice spacing $a_{0}$, such that for any lattice spacing $a<a_{0}$, $\hat{D}$ is always invertible, having no zero mode, along the path in the two parameter space $(m,t)=(-M,0)\to(-M,1)\to(+M,1)\to(+M,0)$. This is a sufficient condition for $[(\mathcal{H}_{\text{cont.}},\gamma(D_{\rm cont.}+m))]$ and $[(\mathcal{H}_{\text{lat.}},\gamma(D_{W}+m))]$ to be in the same equivalence class. We can show that the $\eta(\gamma(D_{W}+M))=0$ for $M>0$ and therefore, we obtain
@@ 5 Comparison with the overlap Dirac index
One possible application is the index on manifolds with boundaries. In the previous sections, we have only considered a periodic square lattice, whose continuum limit is a flat torus. For the massive Dirac operators, it is straightforward to introduce a free boundary condition, to obtain the domain-wall fermion Dirac operator $D_{\rm DW}$ [22, 23, 24]. In [15, 16] we have shown in continuum theory that the $\eta$ invariant of domain-wall Dirac operator is equal to the Atiyah-Patodi-Singer(APS) index of the Dirac operator with a nontrivial boundary. Therefore, it is natural to conjecture that $-\frac{1}{2}\eta(\gamma(D_{\rm DW}))$ at sufficiently small lattice spacings agrees with the APS index. The overlap Dirac operator, in contrast, is no more chiral symmetric when we impose such a boundary condition, since the boundary condition invalidates the Ginsparg-Wilson relation [25].
@@
We also expect that a nontrivial gravitational background can be implemented by a curved domain-wall mass term [27, 28, 29, 30, 31, 31, 32, 33]. It would be interesting to formulate, for instance, the $\hat{A}$ genus of the curved domain-wall on a lattice.